### v1

# A physical random walk for space-fractional diffusion on finite domains

Abstract
We formulate a compounded random walk that is physically well defined on both finite and infinite domains, and samples space-dependent forces throughout jumps. The governing evolution equation for the walk limits to a space-fractional Fokker-Planck equation valid on bounded domains, and recovers the well known superdiffusive space-fractional diffusion equation on infinite domains. This compounded random walk, and its associated fractional Fokker-Planck equation, provides a major advance for modelling space-fractional diffusion through potential fields and on finite domains.
preprint: APS/123-QED
Many physical processes are known to exhibit anomalous diffusion where the mean squared displacement (MSD) scales nonlinearly with time. A standard approach to modeling this is through continuous-time random walks (CTRWs) Montroll and Weiss (1965); Metzler and Klafter (2000). The CTRW describes a random walk where the time between jumps and the jump lengths are stochastic. In the appropriate continuum and diffusion limit, the CTRW master equation can model superdiffusion Metzler and Klafter (2000) where the MSD grows faster than linearly with time. This is of interest for a wide range of physical systems, including: reaction-diffusion equations for evanescent particles Abad et al. (2010), intracellular transport Reverey et al. (2015); Fedotov et al. (2018), random walks with spatial memory Metzler and Klafter (2000); Fedotov et al. (2018), magnetic resonance imaging Magin et al. (2013), electrical propagation in biological tissues Bueno-Orovio et al. (2014), hydrology Benson et al. (2000); Kelly and Meerschaert (2019); Yin et al. (2020); Sun et al. (2020), animal foraging Humphries et al. (2010), laser driven dissipative plasmas Liu and Goree (2008), bacterial swarms Mukherjee et al. (2021) and disordered quantum systems Mildenberger et al. (2007).
Superdiffusion is traditionally generated by jump length distributions with heavy tails and unbounded variance Metzler and Klafter (2000). For superdiffusive CTRWs, the governing equation limits to the space-fractional diffusion equation on the unbounded domain Compte (1996); Metzler and Klafter (2000)
$\dfrac{\partial\rho(x,t)}{\partial t}=-D_{\alpha}(-\nabla^{2})^{\alpha}\rho(x,t),$ (1)
where $\alpha\in(0,1]$ is the fractional exponent, $D_{\alpha}$ is the fractional diffusion coefficient and $\rho(x,t)$ is the probability density function (PDF) of finding the particle at position $x$ at time $t$. Here the fractional Laplacian $-(-\nabla^{2})^{\alpha}$ is defined via a Riesz derivative, which can be expressed in various ways Bayın (2016); Cai and Li (2019); Lischke et al. (2020), with the defining property from the Fourier transform
$\mathcal{F}\{-(-\nabla^{2})^{\alpha}f(x)\}(k)=-|k|^{2\alpha}\mathcal{F}\{f(x)\}(k).$ (2)
The particle motion governed by (1) is characterized by a pseudo superdiffusive MSD, $\lim\limits_{\delta\rightarrow 2\alpha}\langle|x|^{\delta}(t)\rangle\sim t^{\frac{1}{\alpha}}$, and $\delta<2\alpha$ Metzler and Klafter (2000). While the underlying CTRW stochastic process behind (1) is broadly accepted on infinite domains, there are numerous physical problems on finite domains for which there is no general consensus Lischke et al. (2020). Some critical questions arise when considering random walks in finite domains: How can a random walk take its next jump if the jump length exceeds the size of the domain? How can such a non-local random walk incorporate possible effects from spatially varying forces, including potential barriers? These questions are developing areas of research with pertinent examples being stopped Lévy flights Baeumer et al. (2018); Kelly et al. (2019) and subordinated Brownian motion Zoia et al. (2007); Garbaczewski and Stephanovich (2019); Lischke et al. (2020).
Alternatively, the fractional order Laplacian can be defined in the spectral sense through the eigenequation Zoia et al. (2007); Lischke et al. (2020),
$(-\nabla^{2})^{\alpha}v_{n}(x)=\lambda_{n}^{\alpha}v_{n}(x).$ (3)
While the Riesz and spectral fractional Laplacians are equivalent on the infinite domain Zoia et al. (2007); Lischke et al. (2020), the spectral fractional Laplacian remains well defined on finite domains. However, the spectral fractional Laplacian (3) lacks the physical motivation that arises from the diffusion limit of an underlying random walk.
In this Letter, we introduce a compounded random walk whose governing equation gives rise to the space-fractional diffusion equation (1) with the fractional Laplacian defined spectrally (3) in the diffusion limit. The emergence of the spectral Laplacian from the underlying random walk provides a much needed physical interpretation of the space-fractional diffusion equation on bounded domains. Furthermore, this compounded random walk leads to a space-fractional Fokker-Planck equation on both bounded and unbounded domains with proper path sampling of position dependent forces. The random walk interpretation is naturally suited for Monte Carlo simulation methods, which we verify by comparison to both analytical and numerical solutions. Next, we outline the compounded random walk and its governing equation through a CTRW and find spectral representations for the random walker’s probability mass function.
The compounded CTRW model and its solution. Consider a particle, performing a CTRW on a lattice, that waits an exponentially distributed amount of time at site $x_{i}$, with rate $\gamma$, before jumping to a new site $x_{j}$ governed by a jump distribution $\Lambda(x_{j},x_{i})$. The lattice is taken to be one-dimensional with spacing $\Delta x=(b-a)/(m-1)$ on a finite interval $[a,b]$ with $m$ lattice points. We denote $\rho(x_{i},t)$ as the probability of finding the particle at site $x_{i}=a+(i-1)\Delta x$ ($1\leq i\leq m$) at time $t$. Due to probability conservation, $\sum_{j}\Lambda(x_{j},x_{i})=1$. The master equation for the evolution of $\rho(x_{i},t)$ is Montroll and Scher (1973); Angstmann et al. (2015a)
$\dfrac{\partial\rho(x_{i},t)}{\partial t}=\gamma\sum_{j=1}^{m}\Lambda(x_{i},x_{j})\rho(x_{j},t)-\gamma\rho(x_{i},t).$ (4)
Our aim is to find a jump distribution, $\Lambda(x_{i},x_{j})$, for a compounded random walk where a jump consists of a random number of Markovian steps. Figure 1 shows an example of a compounded random walk.
Figure 1: A realization of a single particle performing a compounded CTRW where $\Delta x=1,a=0,b=5$. The top panel shows the lattice site occupancy of the particle where the (orange) horizontal lines show positions and times when the particle is waiting and the (blue) vertical lines show positions sampled by steps during each instantaneous jump. The bottom panels show the $K$ random compounding steps performed by the random walker during jumps at times, $t_{3}$, $t_{4}$ and $t_{8}$.
We begin by defining a nearest neighbor single step distribution. The particle can step one lattice site to the left or right except if the particle attempts to leave the domain ($1\leq i\leq m$), then it remains at its current site. This can be expressed as the single step transition probability,
$\Lambda_{1}(x_{j},x_{i})=r(x_{i})\delta_{j,\min(i+1,m)}+l(x_{i})\delta_{j,\max(i-1,1)}$ (5)
where $\delta_{ji}$ is the Kronecker delta, $r(x_{i})$ is the probability to jump to the right at position $x_{i}$ and $l(x_{i})=1-r(x_{i})$. The probability of a step to the right can be related to a potential, $V(x_{i})$, and an inverse temperature scale, $\beta$, via a Boltzmann weight Henry et al. (2010)
$r(x_{i})=\frac{e^{-\beta V\left(x_{\min(i+1,m)}\right)}}{e^{-\beta V\left(x_{\max(i-1,1)}\right)}+e^{-\beta V\left(x_{\min(i+1,m)}\right)}}.$ (6)
For the compounded jump distribution $\Lambda(x_{i},x_{j})$, in (4), we allow the particle to step a random number of times $K$ in a single jump such that
$\Lambda(x_{j},x_{i})=\sum_{k=1}^{\infty}p_{K}(k)\Lambda_{k}(x_{j},x_{i}),$ (7)
where $p_{K}(k)=\text{Prob}\{K=k\}$. The $k$-step distribution is defined recursively using (5) as
$\Lambda_{k}(x_{j},x_{i})=\sum_{h=1}^{m}\Lambda_{1}(x_{j},x_{h})\Lambda_{k-1}(x_{h},x_{i}),\quad k>1.$ (8)
Substituting the jump distribution (7) into the master equation (4), we obtain
$\begin{split}\frac{\partial\rho(x_{i},t)}{\partial t}=&\gamma\sum_{j=1}^{m}\sum_{k=1}^{\infty}p_{K}(k)\Lambda_{k}(x_{i},x_{j})\rho(x_{j},t)\\ &-\gamma\rho(x_{i},t).\end{split}$ (9)
In order to express the right hand side of (9) spectrally, we first write the eigenequation for the single step distribution,
$\sum_{j=1}^{m}\Lambda_{1}(x_{i},x_{j})X_{n}(x_{j})=(1-\mu_{n})X_{n}(x_{i}),$ (10)
which defines $X_{n}$ as an eigenfunction of the single step distribution with eigenvalue $(1-\mu_{n})$. Then the $k$-step distribution will share the eigenfunctions from (10) such that,
$\sum_{j=1}^{m}\Lambda_{k}(x_{i},x_{j})X_{n}(x_{j})=(1-\mu_{n})^{k}X_{n}(x_{i}).$ (11)
Consequently, the eigenequation for the jump distribution (7) is
$\sum_{j=1}^{m}\Lambda(x_{i},x_{j})X_{n}(x_{j})=G_{K}(1-\mu_{n})X_{n}(x_{i}),$ (12)
where
$G_{K}(z)=\sum_{k=1}^{\infty}p_{K}(k)z^{k},$ (13)
is the probability generating function (PGF), given that $p_{K}(0)=0$.
Now, we use the spectral property of the jump distribution (12) to find the solution of (9) by writing,
$\rho(x_{i},t)=\sum_{n=1}^{m}c_{n}X_{n}(x_{i})e^{-\gamma A_{K,n}t},$ (14)
where for brevity we let $A_{K,n}=1-G_{K}(1-\mu_{n})$ and $c_{n}$ is a coefficient determined by the initial condition $\rho(x_{i},0)$. Differentiating (14) gives
$\frac{\partial\rho(x_{i},t)}{\partial t}=-\gamma\sum_{n=1}^{m}A_{K,n}c_{n}X_{n}(x_{i})e^{-\gamma A_{K,n}t}.$ (15)
Referring back to (4), it follows from the equivalence of the right hand sides of (9) and (15) that any linear spatial operator of the form
$\mathcal{S}\left[f(x_{i})\right]=\gamma\sum_{j=1}^{m}\left[\left(\Lambda(x_{i},x_{j})-\delta_{x_{i},x_{j}}\right)f(x_{j})\right],$ (16)
has an equivalent spectral representation of the form
$\mathcal{S}\left[\rho(x_{i},t)\right]=-\gamma\sum_{n=1}^{m}A_{K,n}c_{n}X_{n}(x_{i})e^{-\gamma A_{K,n}t},$ (17)
where, as above, $A_{K,n}$ can be found from the generating function for the compounded random walk. The validity of (16) and (17) relies on the existence of $m$ distinct eigenvalues. This is ensured as the eigenvalue problem for the single step distribution (10) is equivalent to that for a real tridiagonal matrix with strictly positive values. The jump distribution (12) having $m$ distinct eigenvalues automatically follows as the PGF (13) is monotonically increasing.
Note that this spectral representation is general enough to account for any compounded nearest neighbor random walk in bounded domains and with position dependent potential fields. The master equation for the compounded random walk, (4) and (9), can be written in terms of the linear spatial operator (17) as
$\frac{\partial\rho(x_{i},t)}{\partial t}=\mathcal{S}\left[\rho(x_{i},t)\right].$ (18)
To recover the spectral Laplacian (3) from the discrete space random walk governed by (18), we choose a specific compounding distribution. Using the Sibuya distribution defined as Sibuya (1979); Sibuya and Shimizu (1981)
$p_{K}(k)=\binom{\alpha}{k}(-1)^{k+1},$ (19)
where $0<\alpha\leq 1$, we find the PGF (13) gives
$G_{K}(z)=1-(1-z)^{\alpha}.$ (20)
The compounded random walk with the Sibuya distribution represents a walk where the number of steps within a jump follows a modified Bernoulli process where the probability of stopping on the $k$th step, conditional on not stopping before, is $\alpha/k$. While the Sibuya distribution was used to model power-law waiting times in discrete time random walks Angstmann et al. (2015b), here we use them to model power-law jump lengths.
Substituting the PGF of the Sibuya distribution (20) into the spectral representation of the linear spatial operator for the compounded random walk (17), we obtain
$\mathcal{S}\left[\rho(x_{i},t)\right]=-\gamma\sum_{n=1}^{m}\mu_{n}^{\alpha}c_{n}X_{n}(x_{i})e^{-\gamma\mu_{n}^{\alpha}t}.$ (21)
The linear operator (21) for the unbiased case where $r(x_{i})=1/2$ is a discrete space operator analogous to the spectral fractional Laplacian in (3). Next, we take the diffusion limit of (18) to obtain the fractional diffusion equation (1), where the linear spatial operator (21) reverts to the Riesz derivative (2) on unbounded domains. In addition, we show that the governing equation (18) gives rise to a space-fractional Fokker-Planck equation with a spectrally defined operator on both bounded and unbounded domains.
The diffusion limit. Extending $X_{n}(x_{i})$ to continuous arguments, $X_{n}(x)$, so that we can take a Taylor series, we substitute (5) and (6) into (10). The Taylor series of $X_{n}(x\pm\Delta x)$ about $\Delta x=0$ gives,
$\frac{\Delta x^{2}}{2}\left[2\frac{d}{dx}\left(\beta\frac{dV}{dx}X_{n}\right)+\frac{d^{2}X_{n}}{dx^{2}}\right]+O\left(\Delta x^{4}\right)=-\mu_{n}X_{n}.$ (22)
Using the Fokker-Planck operator,
$\mathcal{L}\left[X_{n}\right]=2\frac{d}{dx}\left(\beta\frac{dV}{dx}X_{n}\right)+\frac{d^{2}X_{n}}{dx^{2}},$ (23)
we can rewrite (22) as
$\frac{\Delta x^{2}}{2}\mathcal{L}\left[X_{n}\right]+O\left(\Delta x^{4}\right)=-\mu_{n}X_{n}.$ (24)
To obtain boundary conditions, we can substitute (5) and (6) into (10) at the boundary points $x_{1}=a$ and $x_{m}=b$ to obtain,
$\Delta x\left[\beta\frac{dV}{dx}X_{n}+\frac{1}{2}\frac{dX_{n}}{dx}\right]\Bigg{|}_{x=a}+O(\Delta x^{2})=-\mu_{n}X_{n}(a),$ (25)
and
$-\Delta x\left[\beta\frac{dV}{dx}X_{n}+\frac{1}{2}\frac{dX_{n}}{dx}\right]\Bigg{|}_{x=b}+O(\Delta x^{2})=-\mu_{n}X_{n}(b).$ (26)
As $\Delta x\rightarrow 0$, we see that the eigenfunctions of the Fokker-Planck operator
$\mathcal{L}\left[\chi_{n}\right]=-\lambda_{n}\chi_{n},$ (27)
must equate to the limit of (24) since $X_{n}\rightarrow\chi_{n}$ in the continuum limit. This equivalence between (24) and (27) gives the relationship
$\lambda_{n}\chi_{n}=\lim_{\Delta x\rightarrow 0}\frac{2\mu_{n}}{\Delta x^{2}}X_{n}.$ (28)
The eigenfunctions $\chi_{n}$ are restricted by the continuum limit of (25) and (26) which impose Robin boundary conditions
$\beta\frac{dV}{dx}\chi_{n}+\frac{1}{2}\frac{d\chi_{n}}{dx}=0,$ (29)
for $x=a$ and $x=b$.
Taking the diffusion limit of (21) and using $\mu_{n}=\Delta x^{2}\lambda_{n}/2$ from (28) as $\Delta x\rightarrow 0$, we obtain
$\mathcal{S}[\rho(x,t)]=-D_{\alpha}\sum_{n=0}^{\infty}\lambda_{n}^{\alpha}c_{n}\chi_{n}(x)e^{-D_{\alpha}\lambda_{n}^{\alpha}t},$ (30)
where the mean waiting time is defined by $\tau=1/\gamma$ and
$D_{\alpha}=\lim\limits_{\Delta x,\tau\rightarrow 0}\frac{\Delta x^{2\alpha}}{2^{\alpha}\tau}.$ (31)
The master equation (18) defined via the linear spatial operator in continuous space (30) becomes,
$\frac{\partial\rho(x,t)}{\partial t}=-D_{\alpha}\sum_{n=0}^{\infty}\lambda_{n}^{\alpha}c_{n}\chi_{n}(x)e^{-D_{\alpha}\lambda_{n}^{\alpha}t}.$ (32)
From (32), we arrive at a space-fractional Fokker-Planck equation,
$\frac{\partial\rho(x,t)}{\partial t}=-D_{\alpha}(-\mathcal{L})^{\alpha}\rho(x,t)$ (33)
where $(-\mathcal{L})^{\alpha}$ is the spectral fractional Fokker-Planck operator. From the eigenequation (27), the spectral fractional Fokker-Planck operator is defined by
$(-\mathcal{L})^{\alpha}\left[\chi_{n}\right]=\lambda_{n}^{\alpha}\chi_{n}.$ (34)
The significance of (33) lies in its connection with an underlying compounded random walk whose $K$ random steps in each jump is Sibuya distributed (19). Note that the fractional Fokker-Planck operator in (33) is fundamentally different to previous formulations that arose from Lévy flights or the Langevin equation on the unbounded domain Fogedby (1994, 1998); Jespersen et al. (1999); Metzler and Klafter (2000). To demonstrate this, we consider the Fourier transform of (33) defined on the unbounded domain
$\frac{\partial\hat{\rho}(k,t)}{\partial t}=-D_{\alpha}(|k|^{2}-ikv)^{\alpha}\hat{\rho}(k,t),$ (35)
where we have set a constant drift, $2\beta dV/dx=v$ in the Fokker-Planck operator (23). In previous models, the fractional Fokker-Planck equation for constant drift $v$ was defined in the Fourier space by Jespersen et al. (1999); Benson et al. (2000); Meerschaert and Tadjeran (2004); Shen et al. (2008)
$\frac{\partial\hat{\rho}(k,t)}{\partial t}=-(D_{\alpha}|k|^{2\alpha}-ikv)\hat{\rho}(k,t).$ (36)
Equation (36) is also known as the fractional advection-dispersion equation Benson et al. (2000); Meerschaert and Tadjeran (2004); Shen et al. (2008); Wang and Barkai (2020). Comparing (35) and (36), it is clear that when $v=0$, the two formulations are equivalent and we obtain the space-fractional diffusion equation (1). This is expected as the Riesz fractional Laplacian is equivalent to the spectral fractional Laplacian for unbounded domains Lischke et al. (2020). However, a clear difference between (35) and (36) is that, in (35) the advection is felt across the entire path of the underlying random walk encapsulated within the term $(|k|^{2}+ikv)^{\alpha}$, as opposed to only accounting for it after landing on a site evident in the term $D_{\alpha}|k|^{2\alpha}+ikv$. For bounded domains, (33) is identical to the fractional diffusion equation defined via the spectral fractional Laplacian with Neumann boundary conditions when $dV/dx=0$ in (29). For a spectral fractional Laplacian formulation with Dirichlet boundary conditions, one needs to reformulate (5) and take the diffusion limit again.
Numerical solutions. The underlying compounded discrete-space random walk for (33) gives a Monte Carlo scheme that can be performed as follows: (1) Partition the interval $[a,b]$ into $m-1$ segments of length $\Delta x=(b-a)/(m-1)$ such that there are $m$ nodes approximating position. (2) Initialize a particle at some initial position $X_{0}=x_{i}$ and set initial time as $t=0$. (3) Draw a random time $\Delta t\sim\text{Exp}(\gamma)$. (4) Draw a Sibuya distributed random variable $K$ Hofert (2011). (5) Given $X=x_{j}$, we will update $X\leftarrow X+\xi\Delta x$ with the following rules: (a) If $1<j<m$, draw $\xi$ with probabilities $\text{Prob}(\xi=1)=r(X)$ and $\text{Prob}(\xi=-1)=1-r(X)$; (b) If $j=1$, then $\text{Prob}(\xi=0)=1-r(X)$ and $\text{Prob}(\xi=1)=r(X)$; (c) If $j=m$, then $\text{Prob}(\xi=0)=r(X)$ and $\text{Prob}(\xi=-1)=1-r(X)$. (6) Repeat step (5) an additional $K-1$ times. (7) Update $t=t+\Delta t$. (8) Repeat steps (3)-(6) until $t\geq T$, where $T$ is the simulation end time. The time discretization of (9) gives a numerical scheme for the solution of (14).
Figure 2: Comparison of solutions between the Monte Carlo simulations (orange), the numerical approximation (purple dots) and the first $1001$ terms of the analytical solution (37) (black line) for the space-fractional diffusion equation on bounded domain $[-1,1]$ with reflecting boundaries. The parameters are $\alpha=0.7$, $N=10^{6}$ particles, $\gamma=10^{2}$, $m=301$ and $T=5$. A subset of data points for the numerical approximation is shown for visual clarity.
Figure 2 shows the Monte Carlo simulation, numerical approximation and the exact spectral solution to (33) with $dV/dx=0$, on the domain $x\in[-1,1]$ with Neumann boundary conditions and initial condition $\rho(x,0)=\delta(x)$. The analytical solution is
$\rho(x,t)=\dfrac{1}{2}+\sum_{n=1}^{\infty}\exp\left(-D_{\alpha}\left(n\pi\right)^{2\alpha}t\right)\cos\left(n\pi x\right).$ (37)
Now, we extend our results to the biased CTRW with a constant forcing term $\beta V^{\prime}(x)=-2$. The analytical solution for the governing equation (33), given initial condition $\rho(x,0)=\delta(x)$, is
$\begin{split}\rho(x,t)=&\dfrac{2e^{4x}}{\sinh(4)}+\sum_{n=1}^{\infty}c_{n}\chi_{n}(x)e^{-D_{\alpha}(4+\omega_{n}^{2})^{\alpha}t},\end{split}$ (38)
where $\omega_{n}=n\pi/2$, $c_{n}=(\omega_{n}\cos\left(\omega_{n}\right)+2\sin\left(\omega_{n}\right))/(4+\omega_{n}^{2})$ and
$\chi_{n}(x)=e^{2x}(\omega_{n}\cos(\omega_{n}(x+1))+2\sin(\omega_{n}(x+1))).$ (39)
Figure 3 shows the correspondence between the Monte Carlo simulation, numerical approximation and the truncated spectral solution (38) for the fractional spectral Fokker-Planck operator with constant bias in (33).
Figure 3: Comparison of solutions between the Monte Carlo simulations (orange), the numerical solution (purple dots) and the first $1001$ terms of the analytical solution (38) (black line) for the space-fractional Fokker-Planck equation on bounded domain $[-1,1]$ with reflecting boundaries, where $\alpha=0.7$, $N=10^{5}$, $\gamma=50$, $L=1$, $m=201$ at multiple times, $T=1$, $2$, $5$ and $10$.
Summary. Our results connect an underlying compounded random walk with the spectral fractional Laplacian in the diffusion limit. This provides a novel and physically valid model to derive the space-fractional diffusion equation on bounded domains. Furthermore, this connection naturally generates a spectral fractional Fokker-Planck equation (33) that is well defined on bounded domains and samples space-dependent forces at every point along a jump. The derivation of the fractional Fokker-Planck equation enables, for the first time, a description of random walkers that experience space-dependent forces over the entire path of a single heavy-tailed jump, improving previous formulations Fogedby (1994, 1998); Jespersen et al. (1999). In doing so, we open an avenue for models with heavy-tailed jumps where the random walk is sampling the environment throughout the jump and is confined to a finite region in space. Moreover, through the PGF, (13), new generalized operators based on experimentally observed microscopic mechanisms can be formulated using (12), (16) and (17) for models of transport, such as intracellular transport Reverey et al. (2015); Fedotov et al. (2018), foraging animals Humphries et al. (2010), and disorderd quantum systems Mildenberger et al. (2007). The governing equations derived through this approach will have convenient spectral solutions and a natural Monte Carlo simulation scheme.
Acknowledgements.
The authors acknowledge financial support by Australian Research Council grant number DP200100345.